OpenStudy (joftheworld):

If a line crosses the y-axis at (0, 1) and has a slope of 4/5, what is the equation of the line?

OpenStudy (anonymous):

Slope intercept form is y = mx + b. m is the slope, and b is the y intercept. We know both of these! Just plug in the values. y = 4/5x + 1.

OpenStudy (anonymous):

Oh, it's in standard form. Get the x to the other side by subtracting. y - 4/5x = 1. Multiply everything by 5 to get rid of the fraction. 5y - 4x = 5.

OpenStudy (joftheworld):

thanks can you help me on a couple more problems ??

OpenStudy (anonymous):

Yep.

OpenStudy (joftheworld):

If a line crosses the y-axis at (0, -4) and has a slope of -2, what is the equation of the line?

OpenStudy (anonymous):

Is the answer supposed to be in standard form?

OpenStudy (anonymous):

y = -2x -4 is the answer in slope intercept. To get standard form, add x to the other side. y + 2x = -4.

OpenStudy (joftheworld):

thanks, one more ?

OpenStudy (anonymous):

Now we do the opposite and turn it into slope intercept by isolating the y. 4x + 5y = -9. Subtract the 4x from the left side. 5y = -9 -4x or 5y = -4x - 9, same thing. Now divide everything by 5 to get y by itself. y = -4/5x - 9/5. We know that y = mx + b where m is the slope and b is the y intercept. The slope is -4/5 and the y intercept is -9/5.